OpenStudy (anonymous):

How do you find the vertex and x intercept of y=4(x-2)^2+1

OpenStudy (anonymous):

@maherman97 Do you know Vertex formula?

OpenStudy (anonymous):

(h,k) is a vertex point. the equation you have is y=a(x-h)^2+k. So just substitute the numbers in. For the x-intercept just set the equation equal to zero.

OpenStudy (anonymous):

@maherman97 @izaikelly has already explained it so neatly and clearly? Please ask if you need further help ?

OpenStudy (anonymous):

I don't understand what to do once you turn the y to zero.

OpenStudy (anonymous):

y = a( x - h )^2 + k. y = 4(x - 2 )^2 + 1

OpenStudy (anonymous):

Now, can you see a = ?

OpenStudy (anonymous):

yesssirr but i dont know what to do with the equation.

OpenStudy (anonymous):

:"find the vertex"

OpenStudy (anonymous):

Im sorry, I really dont understand.

OpenStudy (anonymous):

(h,k) is a vertex

OpenStudy (anonymous):

okay so what do i do to find the x-intercept though?

OpenStudy (anonymous):

4(x - 2 )^2 + 1 = 0

OpenStudy (anonymous):

4(x - 2) ² + 1 = 0 Can you solve it to find x?

OpenStudy (anonymous):

do i do \[4x ^{2}-16x+16\]

OpenStudy (anonymous):

Wait, are you sure your post is correct?

OpenStudy (anonymous):

yeah the problem is \[ y=4(x-2)^{2}+1\]

OpenStudy (anonymous):

you are supposed to graph it too.

OpenStudy (anonymous):

when you set y = 0 4(x - 2) ² + 1 = 0 There's no solution for x

OpenStudy (anonymous):

4(x - 2) ² = -1 Unless you've learned about complex, imaginary solution!

OpenStudy (anonymous):

noo... so what is the x intercept???

OpenStudy (anonymous):

The answer is no x intercept means the graph doesn't cut x axis !

OpenStudy (anonymous):

OHHH, so i just plug in two more points for the graph?

OpenStudy (anonymous):

Yes, if you want to graph it, just plug random values ( choose the one easily to compute)

OpenStudy (anonymous):

okay thanks!